OpenStudy (anonymous):

Show all steps used to solve the problem. One point for steps and one point for solution. Be sure that you check your solution for correctness. B over -3 + 5 =7

OpenStudy (anonymous):

\[\frac{B}{-3}+5=7 \implies \frac{B}{-3}=2\implies B=\frac{-2}{3}\]Do you understand what happened there?

OpenStudy (anonymous):

yes